This is for my Probability and Statistic class,

I need help with this probability problem. We just vaguely just learned conditional and unconditional rules. But I still don't get it.


This is the question and sub-questions involved in the problem. (Given labeled in-class data tables).

Conditional probabilities:
a) Determine the conditional probability that A has visited the (randomly selected) state, given that B has visited it. Is this conditional probability greater than, less than, or the same as the ( unconditional) probability that A has visited the state?
b) Determine the conditional probability that B has visited the (randomly selected) state, given that A has visited it. Is this conditional probability greater than, less than, or the same as the ( unconditional) probability that B has visited the state?
c) Are the events {A has visited the state}, and {B has visited the state} mutually exclusive? Justify your answer.
d) Are the events {A has visited the state}, and {B has visited the state} independent? Justify your answer.
e) Determine the conditional probability that both persons have visited the (randomly selected) state, given that at least one of them has visited it.



Thank you,

Person A and Person B should look through the list and check off which states you have ever visited. (Lets define visited
Answer #1

Let as first define the conditional probability of evente AA and B. - P(A/B) ces ginien the event B, what is the probabilityfrom the table , P(A visted and Bristed2. S P(ARB) = 245 - P(A/B)= P(AMB) _ %0 2 PCBD 5% - % to unconditional probatrility P(Now conditional probability, PROVE PIBIA) = 45 > PCB)-ho e., conditional probability is greater than unconditional probabilitPLAB|Aor or AB) = P(ABIA TO ABIB or ABI AB) = P(ABlA ) +P(ABIB) + PCABIAB) = P(ABNA), PLABOB), P (AB) - PLA) PIB) = P(AB) + P